OpenStudy (anonymous):

How would I find the graph of (-3pi)/5?

OpenStudy (unklerhaukus):

what are the variables ?

OpenStudy (alexwee123):

lol im not really sure :s you're question is kinda vague 

OpenStudy (anonymous):

Thanks man!

OpenStudy (unklerhaukus):

that is theta \(\theta=-\frac{3\pi}{5}\)

OpenStudy (anonymous):

Oh, well that is what I was looking for. Sorry for confusion

OpenStudy (unklerhaukus):

if you were askin for \[x=-\frac {3\pi}{5}\]